OpenStudy (anonymous):

PLZZZ HELP ILL GIVE MEDAL!!! What is the difference? Express your answer in lowest terms. -2/15 - -9/15

OpenStudy (hhmorris98):

7/15

OpenStudy (anonymous):

hows thats the answer? and its supppose to be calculated in lowest terms

OpenStudy (anonymous):

9-2 = 7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx gave u a medal

OpenStudy (anonymous):

i have another one can u help?

OpenStudy (anonymous):

What is the sum? Express your answer in lowest terms. 3/4 + 3/16=

OpenStudy (anonymous):

15/16

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

because 4 times 4 equals 16 so u multiple the numerator by 4

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

plz 1 more medal

OpenStudy (anonymous):

What is the difference? 4/3 - 7/9=

OpenStudy (anonymous):

almost done

OpenStudy (anonymous):

just one more left

OpenStudy (anonymous):

for a medal cause i know its your work

OpenStudy (anonymous):

i gave u a medal

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

dont plz keep it :)

OpenStudy (anonymous):

5/9

OpenStudy (anonymous):

np

OpenStudy (anonymous):

ok, last one

OpenStudy (anonymous):

bruh

OpenStudy (anonymous):

What is the difference? 1/9 - 5/6

OpenStudy (anonymous):

idk sry man

OpenStudy (anonymous):

can u be my fan

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i just need to get this